OpenStudy (royanna01):

if a circle has a circumference of c what is the area of the circle in terms of c

umerlodhi (umerlodhi):

Source: https://answers.yahoo.com/question/index?qid=20140204204712AAnzhWZ A = pi*r² ------------>formula 1 C = 2*pi*r----------->formula 2 solve for r in formula 2, C = 2*pi*r C/(2*pi) = r Substitute r in formula 1, A = pi*r² A = pi*(C/2*pi)² A = C²/4*pi

umerlodhi (umerlodhi):

Medal me if it helped :D